OpenStudy (anonymous):

Can Someone please just help me understand exponents better...Like when you have 2x^3 + 5x^5 and when you subtract, multiply and divide what to do with the exponents...My college alg. teacher went over it a bit but i'm super confused.

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

add or subtract, don't mess with the exponents, leave them alone

OpenStudy (misty1212):

just like \(200+300=500\) you would have \(2x^2+3x^2=5x^2\)

OpenStudy (misty1212):

in your example the \(2x^3\) and \(5x^5\) are not "like terms" so you cannot combine them they are not like terms because the exponents are different

OpenStudy (misty1212):

for example, if \(x\) was \(10\) then \(2x^3=2\times 10^3=2,000\) and \(5x^5=5\times 10^5=500,000\) you can add \[500,000+2,000\] but the 5 and the 2 don't touch, they are in different places

OpenStudy (anonymous):

Ok thanks! What about when you multiply and divide?